# CursorCore: Assist Programming through Aligning Anything

### v2

@@ C H, C C, U H, C, U
Online Submit 27.5 19.7 29.4 23.4
During data collection, we randomly utilize two powerful open-source LLMs: Mistral-Large-Instruct Mistral-AI (2024b) and Deepseek-Coder-V2-Instruct DeepSeek-AI et al. (2024). These models have demonstrated performance comparable to strong closed-source models like GPT-4o across many tasks, and are currently the only two open-source models scoring over 90% on the classic HumanEval benchmark at the start of our experiment. Additionally, they are more cost-effective and offer easier reproducibility than GPT-4o. For Mistral-Large-Instruct, we quantize the model using the GPTQ Frantar et al. (2022) algorithm and deploy it locally with sglang Zheng et al. (2023a) and marlin kernel Frantar et al. (2024) on 4 Nvidia RTX 4090 GPUs. For Deepseek-Coder-V2-Instruct, we use the official API for integration.
Figure 5: The distribution of programming language in the training data.
Figure 6: The distribution of history snippets in the training data.
Figure 7: The distribution of input lengths in the training data.
Figure 8: The distribution of output lengths in the training data.
5.3 Training details
Our models are trained for 2 epochs using the Transformers library Wolf et al. (2020). We enhance memory efficiency and speed with techniques such as Deepspeed ZeRO3 Rajbhandari et al. (2019), ZeRO Offload Ren et al. (2021), FlashAttention2 Dao (2024), and triton kernels Hsu et al. (2024). We calculate the maximum sequence length that can be processed per batch based on the available VRAM. Using the First-Fit Decreasing algorithm Kundu et al. (2024), we pack training samples to ensure that each batch reaches its maximum sequence length, thereby optimizing training speed. The training process employs the Adafactor optimizer Shazeer & Stern (2018) with a learning rate of 5e-5, coupled with a cosine scheduler featuring 15 warm-up steps.
6 Evaluation and Results
6.1 Experimental setup
We conduct the data selection ablation and primary evaluation on our APEval benchmark, and provide results on well-known benchmarks such as Python program synthesis, automated program repair, and instructional code editing, which are detailed in Appendix H. We choose prominent open-source and closed-source LLMs as our baselines. For all benchmarks, we use greedy decoding to generate evaluation results. CursorCore natively supports various inputs in APEval, whereas base and instruction LLMs require additional prompts for effective evaluation. We design few-shot prompts separately for base and instruction models, as detailed in Appendix K.
6.2 Data Selection Ablation
Figure 9: Data Selection Ablation on APEval.
AIprogrammer has the highest data quality
Importance of mixing data with different information types
Our final selection
We combine data from all sources for training. Since our focus is on Python, and training on multilingual data leads to a decrease in APEval scores, we use only the Python part of the Git Commit and Online Submit datasets. As a result, we get CursorCore series models.
@@ 6.3 Evaluation results on APEval
In Table 4, we present the results of evaluating CursorCore series models and other LLMs on APEval. It includes both the average results and the results across four different types of information within the benchmark, each item in the table is the score resulting from running the base tests and extra tests. We also report the evaluation results of other well-known models, which can be found in Appendix I.
@@
For most model series, instruction-tuned models outperform their corresponding base models, as instruction fine-tuning generally enhances model capabilities Ouyang et al. (2022); Longpre et al. (2023). The only exception observed in our experiments is the latest model, Qwen2.5-Coder. Its base model achieves a very high score, while the instruction-tuned model performes worse. We attribute the base model’s high performance to its extensive pre-training, which involved significantly more tokens than previous models Hui et al. (2024). This training on a wide range of high-quality data grants it strong generalization abilities, enabling it to effectively handle the newly defined APEval task format. In contrast, the instruction-tuned model is not specifically aligned with this task, leading to a decrease in its APEval score. This highlights the challenges of aligning models with numerous diverse tasks, especially small models.
@@
Historically, open-source models significantly lag behind closed-source models, like those in the GPT series, leading to a preference for closed-source models in synthetic data generation and other applications Taori et al. (2023); Xu et al. (2023). However, with the continuous advancement of open-source LLMs, increasingly powerful models have emerged. On APEval, the best open-source models—such as Qwen2.5-72B-Instruct, Mistral-Large-Instruct, and Deepseek-Coder-V2-Instruct—demonstrate performance that closely approaches that of the leading GPT series model, GPT-4o. This indicates that the performance gap between open-source and closed-source LLMs has considerably narrowed, encouraging the development of more interesting applications based on open-source LLMs. Despite this progress, GPT-4o remains more comprehensive than open-source LLMs. It utilizes H far more effectively than any other model, demonstrating its strong capability to process and align with various types of information. This is an area where open-source LLMs still need to improve.
@@ A.1 AI-Assisted Programming
AI-assisted programming has a long history, encompassing various tasks such as clone detection Lu et al. (2021), code summarization Sun et al. (2024), program synthesis Chen et al. (2021); Austin et al. (2021), automatic program repair Gulwani et al. (2016), code editing Wei et al. (2023a), and code optimization Shypula et al. (2024). These tasks attempt to incorporate a wide range of information into their processes, such as historical edits Gupta et al. (2023); Zhang et al. (2022) and user instructions Cassano et al. (2023b). In the past, however, they were typically addressed by custom-built models, which were difficult to scale across different tasks and types of information. With the rise of LLMs, AI-assisted programming increasingly leverages LLMs to handle multiple types of tasks simultaneously. Numerous high-quality open-source and closed-source products, such as Continue Continue-Dev (2024), Aider Paul-Gauthier (2024), Copilot Github-Copilot (2022) and Cursor Cursor-AI (2023), are based on this approach.
@@ A.2 Code Models
Recently, LLMs have attracted significant attention in the research community for their impact on enhancing various aspects of code intelligence. Open-source code LLMs like CodeLlama Rozière et al. (2023); Touvron et al. (2023), Deepseek-Coder Guo et al. (2024a); DeepSeek-AI et al. (2024), StarCoder Li et al. (2023); Lozhkov et al. (2024), Codegemma Team et al. (2024), Codestral Mistral-AI (2024a), Codegeex Zheng et al. (2023b), Yi-Coder AI et al. (2024), and Qwen-Coder Hui et al. (2024) have made substantial contributions by utilizing large code corpora during training. Some models, such as WizardCoder Luo et al. (2024b), OctoCoder Muennighoff et al. (2024), CodeLlama-Instruct, Deepseek-Coder-Instruct, MagiCoder Wei et al. (2023b), Yi-Coder-Chat, and Qwen-Coder-Instruct, have been fine-tuned using instruction data collected through methods like Self-Instruct Wang et al. (2023); Taori et al. (2023), Evol-Instruct, and OSS-Instruct. These models are specifically trained on code-related instructions, improving their ability to follow coding instructions. They have made significant breakthroughs in tasks like code completion and editing.
@@ A.3 Code Benchmarks
HumanEval Chen et al. (2021) is one of the most well-known benchmarks in the code domain, featuring several variants that extend it to different programming languages, extra tests, and broader application scenarios. Other notable benchmarks include MBPP Austin et al. (2021) for program synthesis, DS1000 Lai et al. (2022) for data science tasks, SWE-Bench Jimenez et al. (2024) for real-world software engineering problems, and CanItEdit / CodeEditorBench Cassano et al. (2023b); Guo et al. (2024b) for code editing. Additionally, LiveCodeBench Jain et al. (2024) focuses on contamination-free evaluations, while Bigcodebench Zhuo et al. (2024) and Naturecodebench Zhang et al. (2024b) provide comprehensive program synthesis assessments. CRUXEval Gu et al. (2024) targets reasoning, CrossCodeEval Ding et al. (2023) focuses on repository-level code completion, and Needle in the code Hui et al. (2024) is designed for long-context evaluations.
@@ Appendix B Code modification representation
As discussed in Section 2.3, there are various ways to represent code modifications. Many previous works have explored techniques for instruction-based code editing Wei et al. (2023a); Muennighoff et al. (2024); Paul-Gauthier (2024); Sweep-AI (2024). We build upon these works with the following formats, as shown in Figure 12:
@@
We use the entire code, allows for a straightforward representation of the modifications. However, when only small parts of the code are changed, this method leads to redundancy, especially for long code files. Certain mitigation can be achieved through technologies such as retrieval-based speculative decoding Yang et al. (2023); He et al. (2024).
@@
Recent research has pointed out correlations and scaling laws between model input and output length, as well as performance OpenAI (2024); Snell et al. (2024). Our results align with these findings. As the length increases, performance improves consistently across LC, SR, and WF. UD performs poorly in both token usage and performance, likely because it contains redundant information, such as both line numbers and content for the modified sections, where only one would suffice. This redundancy reduces the format’s efficiency compared to the other three formats.
@@ Appendix C Details regarding the collection process of APEval
In our benchmark collection process, we initially annotated one programming process for each task. For some tasks, the annotators consulted the programming assistant; for others, they did not. Similarly, some tasks involved complex editing histories, while others did not. Upon reviewing the data, we found that for certain tasks, it was nearly impossible to collect realistic programming processes containing specific types of information. For example, Some tasks are straightforward and can be completed with just a few lines of code. Programmers who have undergone basic training can write these solutions quickly without needing to consult an assistant or repeatedly revise their code. Conversely, some tasks may involve calling specific libraries or algorithms that most annotators are unfamiliar with, leading them to rely on the programming assistant. It would be unrealistic and counterproductive to instruct annotators to ”always consult the AI” or ”edit your code repeatedly,” as this would deviate from real-world scenarios and undermine our intention to use human-annotated data. Considering these reasons, we did not collect programming traces for the entire test set. While we still hope that the number of samples of four different combinations is at least balanced. At this stage, the number of samples for combinations involving all four data types was relatively similar. So we asked annotators to label additional programming process traces for combinations with fewer samples and collected the corresponding traces. Meanwhile, for combinations with slightly more samples, we discarded some of their traces. Through this process, we established our final benchmark. Simplified examples of the annotated data is illustrated in Figure 13.
@@ Appendix D Additional details about Programming-Instruct
In our code editing records, we place no limits on the granularity or number of edits. Changes between two code versions may involve anything from a single character to multiple extensive modifications. However, data collected from various sources may be compressed, resulting in incomplete records. This compression can lead to a higher proportion of large-scale edits, particularly in Git Commit data. To address this issue, we propose a decomposition strategy: when there are multiple changes between versions, we break them down into single-step modifications, with the steps ordered randomly. For Git Commit data, we apply this decomposition strategy with a 90% probability, while for AIprogrammer and Online Submit data, we apply it with a 50% probability.
@@
In addition to generating H and U, as discussed in Section 4.2, we also simulate the programmer’s specification of the target area and model interactions in a chat-style format. The target modification area is created using a random algorithm, as described in Appendix E, while the chat-style interaction is generated using LLMs which is similar to the generation of instructions. Prompts used for it are provided in Appendix L.
Appendix E Target area representation
@@
We evaluate CursorCore-DS-1.3B on APEval both with and without location information to assess its impact on performance. The results in Figure 14 show that including location information has minimal effect, likely because most APEval examples are relatively short, enabling LLMs to easily infer modification locations, much like humans do without a cursor. Previous works, such as those on automated program repair Zhang et al. (2024a), have emphasized the importance of identifying the modification location. We believe this emphasis stems from traditional code completion and insertion paradigms, as well as the natural alignment of specifying modification points with human thought processes. However, with the advancement of LLMs, the benefit of providing location information diminishes when generating code at the function or file level. This may need further exploration in longer contexts, such as repository-level editing tasks.
Appendix F Discussion about thought process
@@
Incorporating reasoning processes in prompts has been shown to improve model performance, as demonstrated in various works like CoT Wei et al. (2022) and ReACT Yao et al. (2023). Some studies have even integrated these processes into the training phase to further enhance effectiveness Zelikman et al. (2022). In this work, we also explore a self-taught approach, where we prompt LLMs to reverse-generate the reasoning process from outputs and incorporate them into the model’s output during training. Our model and data setup follow the same configuration as described in Appendix B to enable quick experiments. The evaluation results are shown in Figure 15.
@@
Appendix G Conversation retrieval for Assistant-Conversation
Figure 16: Performance of models using different sliding window sizes on APEval.
@@ Appendix G Conversation retrieval for Assistant-Conversation
To address this, it is essential to compress the editing history or retrieve only the relevant portions. Similar to how many conversation retrieval techniques, such as memory modules Packer et al. (2023), prompt compression Jiang et al. (2023) and query rewriting Ye et al. (2023), are used to manage dialogues for chatbots, these methods can be adapted for handling code editing records. In this work, we explore a basic approach, sliding window, to investigate possible solutions. When the number of historical editing records surpasses a predefined threshold, the model automatically discards the oldest entries.
@@
Appendix H Evaluation results of other benchmarks
We also evaluate CursorCore on other well-known benchmarks. We use HumanEval+ and MBPP+ Liu et al. (2023) to evaluate Python program synthesis, CanItEdit Cassano et al. (2023b) for instructional code editing, and the Python subset of HumanEvalFix from OctoPack Muennighoff et al. (2024) for automated program repair. All benchmarks are based on their latest versions, and HumanEvalFix uses the test-based repair version as described in the original paper. To generate results, we consistently use vLLM Kwon et al. (2023) due to its versatility and support for customized conversation formats. Evaluations are conducted within each benchmark’s execution environment.
  • Chat: Similar to the chat format of ChatGPT Ouyang et al. (2022), we wrap the query before passing it to the model, which returns a response in a chat style. The final result is obtained after post-processing.
  • Inline: Similar to Copilot Inline Chat Github-Copilot (2022) and Cursor Command K Cursor-AI (2023) scenarios, corresponding to the combination of C and U in Assistant-Conversation. Compared to the Chat mode, it is more tightly integrated with the IDE and returns less additional content.
  • Tab: Similar to the use case of Copilot++ Cursor-AI (2023), it is the most automated of all scenarios. We provide only the C to the model. For instructional code editing and automated code repair, no explicit instructions are passed.
@@ Appendix H Evaluation results of other benchmarks
Appendix I Additional evaluation results on APEval

### v3

@@ C H, C C, U H, C, U
Online Judge Submission 27.5 19.7 29.4 23.4
@@
6 Evaluation and Results
6.1 Experimental setup
We conduct the data selection ablation and primary evaluation on our APEval benchmark, and provide results on well-known benchmarks such as Python program synthesis, automated program repair, and instructional code editing, which are detailed in Appendix J. We choose prominent open-source and closed-source LLMs as our baselines. For all benchmarks, we use greedy decoding to generate evaluation results. CursorCore natively supports various inputs in APEval, whereas base and instruction LLMs require additional prompts for effective evaluation. We design few-shot prompts separately for base and instruction models, as detailed in Appendix N. Data selection ablation can be found in Appendix H.
@@ 6.2 Evaluation results on APEval
In Table 4, we present the results of evaluating CursorCore series models and other LLMs on the Python version of APEval. The results for multilingual versions can be found in Appendix L. It includes both the average results and the results across four different types of information within the benchmark, each item in the table is the score resulting from running the base tests and extra tests. We also report the evaluation results of other well-known models, which can be found in Appendix K.
@@
For most model series, instruction-tuned models outperform their corresponding base models, as instruction fine-tuning generally enhances model capabilities (Ouyang et al., 2022; Longpre et al., 2023). The only exception observed in our experiments is the latest model, Qwen2.5-Coder. Its base model achieves a very high score, while the instruction-tuned model performes worse. We attribute the base model’s high performance to its extensive pre-training, which involved significantly more tokens than previous models (Hui et al., 2024). This training on a wide range of high-quality data grants it strong generalization abilities, enabling it to effectively handle the newly defined APEval task format. In contrast, the instruction-tuned model is not specifically aligned with this task, leading to a decrease in its APEval score. This highlights the challenges of aligning models with numerous diverse tasks, especially small models.
@@
Historically, open-source models significantly lag behind closed-source models, like those in the GPT series, leading to a preference for closed-source models in synthetic data generation and other applications (Taori et al., 2023; Xu et al., 2023). However, with the continuous advancement of open-source LLMs, increasingly powerful models have emerged. On APEval, the best open-source models—such as Qwen2.5-72B-Instruct, Mistral-Large-Instruct, and Deepseek-Coder-V2-Instruct—demonstrate performance that closely approaches that of the leading GPT series model, GPT-4o. This indicates that the performance gap between open-source and closed-source LLMs has considerably narrowed, encouraging the development of more interesting applications based on open-source LLMs. Despite this progress, GPT-4o remains more comprehensive than open-source LLMs. It utilizes H far more effectively than any other model, demonstrating its strong capability to process and align with various types of information. This is an area where open-source LLMs still need to improve.
@@ 7 Conclusion
Acknowledgments
This research was partially supported by grants from the Joint Research Project of the Science and Technology Innovation Community in Yangtze River Delta (No. 2023CSJZN0200), the National Natural Science Foundation of China (62337001), the Key Technologies R & D Program of Anhui Province (No. 202423k09020039) and the Fundamental Research Funds for the Central Universities.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ A.1 AI-Assisted Programming
AI-assisted programming has a long history, encompassing various tasks such as clone detection (Lu et al., 2021), knowledge tracing (Liu et al., 2019; Li et al., 2022; Gao et al., 2025), data mining (Wang et al., 2023a; Zhang et al., 2024c), code retrieval (Li et al., 2024b, c, 2025), code summarization (Jiang et al., 2025; Sun et al., 2024), program synthesis (Chen et al., 2021; Austin et al., 2021), automatic program repair (Gulwani et al., 2016; Zhao et al., 2024), code editing (Wei et al., 2023a), and code optimization (Shypula et al., 2024). These tasks attempt to incorporate a wide range of information into their processes, such as historical edits (Gupta et al., 2023; Zhang et al., 2022) and user instructions (Cassano et al., 2023b). In the past, however, they were typically addressed by custom-built models, which were difficult to scale across different tasks and types of information. With the rise of LLMs, AI-assisted programming increasingly leverages LLMs to handle multiple types of tasks simultaneously. Numerous high-quality open-source and closed-source products, such as Continue (Continue-Dev, 2024), Aider (Paul-Gauthier, 2024), Copilot (Github-Copilot, 2022) and Cursor (Cursor-AI, 2023), are based on this approach.
@@ A.2 Code Models
Recently, LLMs have attracted significant attention in the research community for their impact on enhancing various aspects of code intelligence. Open-source code LLMs like CodeLlama (Rozière et al., 2023; Touvron et al., 2023), Deepseek-Coder (Guo et al., 2024a; DeepSeek-AI et al., 2024), StarCoder (Li et al., 2023; Lozhkov et al., 2024), Codegemma (Team et al., 2024), Codestral (Mistral-AI, 2024a), Codegeex (Zheng et al., 2023b), Yi-Coder (AI et al., 2024), and Qwen-Coder (Hui et al., 2024) have made substantial contributions by utilizing large code corpora during training. Some models, such as WizardCoder (Luo et al., 2024b), OctoCoder (Muennighoff et al., 2024), CodeLlama-Instruct, Deepseek-Coder-Instruct, MagiCoder (Wei et al., 2023b), Yi-Coder-Chat, and Qwen-Coder-Instruct, have been fine-tuned using instruction data collected through methods like Self-Instruct (Wang et al., 2023b; Taori et al., 2023), Evol-Instruct, and OSS-Instruct. These models are specifically trained on code-related instructions, improving their ability to follow coding instructions. They have made significant breakthroughs in tasks like code completion and editing.
@@ A.3 Code Benchmarks
HumanEval (Chen et al., 2021) is one of the most well-known benchmarks in the code domain, featuring several variants that extend it to different programming languages, extra tests, and broader application scenarios. Other notable benchmarks include MBPP (Austin et al., 2021) for program synthesis, DS1000 (Lai et al., 2022) for data science tasks, SWE-Bench (Jimenez et al., 2024) for real-world software engineering problems, and CanItEdit / CodeEditorBench (Cassano et al., 2023b; Guo et al., 2024b) for code editing. Additionally, LiveCodeBench (Jain et al., 2024) focuses on contamination-free evaluations, while ClassEval(Du et al., 2023), Bigcodebench (Zhuo et al., 2024) and Naturecodebench (Zhang et al., 2024b) provide comprehensive program synthesis assessments. CRUXEval (Gu et al., 2024) targets reasoning, CrossCodeEval (Ding et al., 2023) focuses on repository-level code completion, and Needle in the code (Hui et al., 2024) is designed for long-context evaluations.
@@ Appendix B Code modification representation
As discussed in Section 2.3, there are various ways to represent code modifications. Many previous works have explored techniques for instruction-based code editing (Wei et al., 2023a; Muennighoff et al., 2024; Paul-Gauthier, 2024; Sweep-AI, 2024). We build upon these works with the following formats, as shown in Figure 11:
@@
We use the entire code, allows for a straightforward representation of the modifications. However, when only small parts of the code are changed, this method leads to redundancy, especially for long code files. Certain mitigation can be achieved through technologies such as retrieval-based speculative decoding (Yang et al., 2023; He et al., 2024).
@@
Recent research has pointed out correlations and scaling laws between model input and output length, as well as performance (OpenAI, 2024; Snell et al., 2024). Our results align with these findings. As the length increases, performance improves consistently across LC, SR, and WF. UD performs poorly in both token usage and performance, likely because it contains redundant information, such as both line numbers and content for the modified sections, where only one would suffice. This redundancy reduces the format’s efficiency compared to the other three formats.
@@ Appendix C Details regarding the collection process of APEval
In our benchmark collection process, we initially annotated one programming process for each task. For some tasks, the annotators consulted the programming assistant; for others, they did not. Similarly, some tasks involved complex editing histories, while others did not. Upon reviewing the data, we found that for certain tasks, it was nearly impossible to collect realistic programming processes containing specific types of information. For example, Some tasks are straightforward and can be completed with just a few lines of code. Programmers who have undergone basic training can write these solutions quickly without needing to consult an assistant or repeatedly revise their code. Conversely, some tasks may involve calling specific libraries or algorithms that most annotators are unfamiliar with, leading them to rely on the programming assistant. It would be unrealistic and counterproductive to instruct annotators to ”always consult the AI” or ”edit your code repeatedly,” as this would deviate from real-world scenarios and undermine our intention to use human-annotated data. Considering these reasons, we did not collect programming traces for the entire test set. While we still hope that the number of samples of four different combinations is at least balanced. At this stage, the number of samples for combinations involving all four data types was relatively similar. So we asked annotators to label additional programming process traces for combinations with fewer samples and collected the corresponding traces. Meanwhile, for combinations with slightly more samples, we discarded some of their traces. Subsequently, we manually translated them into different programming languages. Through this process, we established our final benchmark. Simplified examples of the annotated data is illustrated in Figure 12.
@@ Appendix D Additional details about Programming-Instruct
In our code editing records, we place no limits on the granularity or number of edits. Changes between two code versions may involve anything from a single character to multiple extensive modifications. However, data collected from various sources may be compressed, resulting in incomplete records. This compression can lead to a higher proportion of large-scale edits, particularly in Git Commit data. To address this issue, we propose a decomposition strategy: when there are multiple changes between versions, we break them down into single-step modifications, with the steps ordered randomly. For Git Commit data, we apply this decomposition strategy with a 90% probability, while for AI Programmer and Online Judge Submission data, we apply it with a 50% probability.
@@
In addition to generating H and U, as discussed in Section 4.2, we also simulate the programmer’s specification of the target area and model interactions in a chat-style format. The target modification area is created using a random algorithm, as described in Appendix F, while the chat-style interaction is generated using LLMs which is similar to the generation of instructions. Prompts used for it are provided in Appendix O.
Appendix E Training details
Our models are trained for 2 epochs using the Transformers library (Wolf et al., 2020). We enhance memory efficiency and speed with techniques such as Deepspeed ZeRO3 (Rajbhandari et al., 2019), ZeRO Offload (Ren et al., 2021), FlashAttention2 (Dao, 2024), and triton kernels (Hsu et al., 2024). We calculate the maximum sequence length that can be processed per batch based on the available VRAM. Using the First-Fit Decreasing algorithm (Kundu et al., 2024), we pack training samples to ensure that each batch reaches its maximum sequence length, thereby optimizing training speed. The training process employs the Adafactor optimizer (Shazeer & Stern, 2018) with a learning rate of 5e-5, coupled with a cosine scheduler featuring 15 warm-up steps.
Appendix F Target area representation
@@
We evaluate CursorCore-DS-1.3B on APEval both with and without location information to assess its impact on performance. The results in Figure 13 show that including location information has minimal effect, likely because most APEval examples are relatively short, enabling LLMs to easily infer modification locations, much like humans do without a cursor. Previous works, such as those on automated program repair (Zhang et al., 2024a), have emphasized the importance of identifying the modification location. We believe this emphasis stems from traditional code completion and insertion paradigms, as well as the natural alignment of specifying modification points with human thought processes. However, with the advancement of LLMs, the benefit of providing location information diminishes when generating code at the function or file level. This may need further exploration in longer contexts, such as repository-level editing tasks.
Appendix G Discussion about thought process
@@
Incorporating reasoning processes in prompts has been shown to improve model performance, as demonstrated in various works like CoT (Wei et al., 2022) and ReACT (Yao et al., 2023). Some studies have even integrated these processes into the training phase to further enhance effectiveness (Zelikman et al., 2022). In this work, we also explore a self-taught approach, where we prompt LLMs to reverse-generate the reasoning process from outputs and incorporate them into the model’s output during training. Our model and data setup follow the same configuration as described in Appendix B to enable quick experiments. The evaluation results are shown in Figure 14.
@@
Appendix H Data Selection Ablation
Figure 15: Data Selection Ablation on APEval.
AI Programmer has the highest data quality
Importance of mixing data with different information types
Our final selection
We combine data from all sources for training. Since current research on Code LLMs primarily focuses on performance in Python, and training on multilingual data leads to a slight decrease in APEval scores, we use only the Python part of the Git Commit and Online Judge Submission datasets. As a result, we get CursorCore series models.
Appendix I Conversation retrieval for Assistant-Conversation
Figure 16: Performance of models using different sliding window sizes evaluated on APEval.
@@ Appendix I Conversation retrieval for Assistant-Conversation
To address this, it is essential to compress the editing history or retrieve only the relevant portions. Similar to how many conversation retrieval techniques, such as memory modules (Packer et al., 2023), prompt compression (Jiang et al., 2023) and query rewriting (Ye et al., 2023), are used to manage dialogues for chatbots, these methods can be adapted for handling code editing records. In this work, we explore a basic approach, sliding window, to investigate possible solutions. When the number of historical editing records surpasses a predefined threshold, the model automatically discards the oldest entries.
@@
Appendix J Evaluation results of other benchmarks
@@
We also evaluate CursorCore on other well-known benchmarks. We use HumanEval+ and MBPP+ (Liu et al., 2023) to evaluate Python program synthesis, CanItEdit (Cassano et al., 2023b) for instructional code editing, and the Python subset of HumanEvalFix from OctoPack (Muennighoff et al., 2024) for automated program repair. All benchmarks are based on their latest versions, and HumanEvalFix uses the test-based repair version as described in the original paper. To generate results, we consistently use vLLM (Kwon et al., 2023) due to its versatility and support for customized conversation formats. Evaluations are conducted within each benchmark’s execution environment.
  • Chat: Similar to the chat format of ChatGPT (Ouyang et al., 2022), we wrap the query before passing it to the model, which returns a response in a chat style. The final result is obtained after post-processing.
  • Inline: Similar to Copilot Inline Chat (Github-Copilot, 2022) and Cursor Command K (Cursor-AI, 2023) scenarios, corresponding to the combination of C and U in Assistant-Conversation. Compared to the Chat mode, it is more tightly integrated with the IDE and returns less additional content.
  • Tab: Similar to the use case of Copilot++ (Cursor-AI, 2023), it is the most automated of all scenarios. We provide only the C to the model. For instructional code editing and automated code repair, no explicit instructions are passed.
Table 6: Evaluation results on Zeta, DS1000 and ClassEval.
Model Zeta DS1000 ClassEval
DS-Coder-1.3B-Base 18.2 16.2 13.0
DS-Coder-1.3B-Inst 42.4 20.7 13.0
CursorCore-DS-1.3B 45.5 21.2 17.0
To further evaluate the ability of CursorCore to leverage historical information for editing and its applicability to more general software engineering tasks, we additionally conduct experiments on Zeta (Zed-Industries, 2025), DS1000 (Lai et al., 2022), and ClassEval (Du et al., 2023), as shown in Table 6. For Zeta, we report the average accuracy across all evaluated samples, with correctness judged by GPT-4o based on the associated assertion text. For DS1000 and ClassEval, we choose to use the Inline and Tab modes, as they most closely resemble the original formats of them. We report the average score across all samples, using the subset of ClassEval that evaluates class-level generation. All generations are produced under greedy decoding. These results collectively demonstrate the strong effectiveness of CursorCore.
Appendix K Additional evaluation results on APEval
@@
Appendix L Multilingual evaluation results on APEval
We report the evaluation results on multilingual versions of APEval, as shown in Tables 8, 9, 10, 11, 12 and 13. CursorCore series achieve state-of-the-art performance across all languages, strongly demonstrating the effectiveness of our approach.